Graph the following. For each one, identify the vertex, and the axis of symmetry.
Describe the transformations.
a. y = x 2 + 9
b. y = -2(x - 1) 2
c. y = -2(x - 1) 2 + 2
d. y = 1 / 2 (x + 3) 2 - 1

Answer by ewatrrr(24785)


Hi
a. y = x^2 + 9                V(0,9)  axis of symmetry:  x = 0
b. y = -2(x - 1)^2            V(1,0)  axis of symmetry:  x = 1
c. y = -2(x - 1)^2 + 2        V(1,2)  axis of symmetry:  x = 1
d. y = 1/2 (x + 3)^2 - 1      V(-3,-1)  axis of symmetry:  x =-3



Wish You the Best in your Studies.